what is the slope of the line given by the equation below ?
y = 6x

Answer by stanbon(75887)

what is the slope of the line given by the equation below ?
y = 6x
----
The slope is the amount y changes when x changes by 1.
If x = 0, y = 6*0 = 0
If x = 1, y = 6*1 = 6
------
So the slope = 6
===============
Cheers,
Stan H.